Water from a vertical pipe emerges as a 18-cm-diam
Answer #1

Velocity at the bottom will be

v_{2} = \sqrt{v_{1}^{2}+2gh} = \sqrt{(2m/s)^{2}+2\times 9.8 m/s^{2}\times 7.5 m } = 12.29 m/s

We know from equation of continuity

a_{1}v_{1} = a_{2}v_{2}

\pi \left (\frac{D_{1}}{2} \right )^{2}v_{1} = \pi \left (\frac{D_{2}}{2} \right )^{2}v_{2}

D_{1}^{2}v_{1} = D_{2}^{2}v_{2}

D_{2}= \sqrt{\frac{D_{1}^{2}v_{1}}{v_{2} }}

D_{2}= \sqrt{\frac{(18\times 10^{-2}m)^{2}\times 2.0 m/s}{12.29 m/s }}= 7.26\times 10^{-2} m = 7.26 cm\approx 7.3 cm